# SimpleToM: Exposing the Gap between Explicit ToM Inference and Implicit ToM Application in LLMs

### v1

@@ Abstract
While prior work has explored whether large language models (LLMs) possess a “theory of mind” (ToM) - the ability to attribute mental states to oneself and others - there has been little work testing whether LLMs can implicitly apply such knowledge to predict behavior, or to judge whether an observed behavior is rational. Such skills are critical for appropriate interaction in social environments. Our approach to study such capabilities is to create a new dataset, called SimpleToM, containing concise, diverse stories (e.g., “The can of Pringles has moldy chips in it. Mary picks up the can in the supermarket and walks to the cashier.”), each with three questions that test different degrees of ToM reasoning, asking models to predict (a) mental state (“Is Mary aware of the mold?”), (b) behavior (“Will Mary pay for the chips or report the mold?”), and (c) judgment (“Mary paid for the chips. Was that reasonable?”). To our knowledge, SimpleToM is the first dataset to systematically explore downstream reasoning requiring knowledge of mental states in realistic scenarios. Our experimental results are intriguing: While most models can reliably predict mental state on our dataset (a), they often fail to correctly predict the behavior (b), and fare even worse at judging whether given behaviors are reasonable (c), despite being correctly aware of the protagonist’s mental state should make such secondary predictions obvious. We further show that we can help models do better at (b) and (c) via interventions such as reminding the model of its earlier mental state answer and mental-state-specific chain-of-thought prompting, raising the action prediction accuracies (e.g., from 49.5% to 93.5% for GPT-4o) and judgment accuracies (e.g., from 15.3% to 94.7% in GPT-4o). However, while this shows that models can be coaxed to perform well, it requires task-specific interventions, and the natural model performances remain low, a cautionary tale for LLM deployment. SimpleToM thus breaks new ground in probing real-world ToM reasoning, and reveals surprising, new insights about current model capabilities. We hope the dataset enables further exploration by the community into this critical area of model behavior.
@@ 1 Introduction
People infer what others know, anticipate their actions, and expect them to choose cost-minimizing behaviors (Gergely & Csibra, 2003; Liu & Spelke, 2017). Performing such social reasoning involves attributing mental states to oneself and others, an ability widely known as Theory of Mind (ToM) (Premack & Woodruff, 1978). ToM has been extensively studied in psychology in a range of scenarios, for instance, studies of manipulation, secrecy, deception, lying, and misleading behavior (Doherty, 2008). Examples of classical tests in developmental psychology include the unexpected transfer false belief task, e.g., the Sally-Anne task (Baron-Cohen et al., 1985), and the unexpected contents false belief task, e.g., the Smarties task (Perner et al., 1987). Given the increasing use of LLMs in human interactions and as decision-making agents within complex, human-centered environments, it is crucial to assess their ToM capabilities.
@@
Current datasets for assessing ToM in LLMs are limited by their reliance on the classical Sally-Anne task or templated variants of it (Le et al., 2019; Nematzadeh et al., 2018; Wu et al., 2023; Xu et al., 2024). These datasets have significant shortcomings: (i) limited diversity in how information asymmetry arises (see related work in Section 6 for examples across existing datasets), (ii) explicit use of percept and mentalizing verbs like “sees” and “thinks” which serve as trigger words for models to realize that these are important aspects, removing the need for implicit commonsense inferences about relevant percepts or beliefs, and (iii) limited exploration of applied ToM, such as judgment of behavior which requires implicit reasoning about mental state.
@@
The goal of SimpleToM is to go beyond the focus on mental state inference, to also assess how well models can predict others’ behavior based on that understanding, and make judgments of appropriateness of behavior. Each story in SimpleToM is paired with three types of questions targeting these abilities (Figure 1), with a total of 1147 stories and 3441 questions in daily life settings. Our analysis using SimpleToM exposes a significant gap in model performance between explicit and applied ToM questions (Lee et al., 2024). We find that frontier models perform well on explicit ToM questions (directly querying for information about “mental state”, i.e., information awareness). However, this success does not extend to applied ToM (“behavior” and “judgment” questions). Scenario-specific analysis highlights the need for diverse testing scenarios to rigorously assess neural ToM and indicates areas for improvement even in strong models like o1-preview. Our experiments with various inference-time interventions show that the performance of applied ToM improves with increased guidance. However, the need for highly specific guidance and the high inference cost (e.g., CoT using hundreds of tokens) required for models to reason through simple two-sentence stories in SimpleToM, suggests that the inherent ToM reasoning abilities of advanced LLMs still have significant room for improvement. The results showing that frontier models still lack the ability to independently apply ToM skills in downstream tasks—such as anticipating others’ behavior and making judgments—call for caution when using them in real-world applications.
Our contributions and findings are:
@@
  • A new dataset, SimpleToM, for testing core abilities of LLMs in explicit and applied ToM.
@@
  • Current frontier models have decoupled capabilities between predicting someone’s information awareness in a situation (explicit ToM, which they excel at), and utilizing it to predict and judge someone’s behavior (applied ToM, which they perform poorly at).
  • Models can be explicitly guided to perform applied ToM, e.g., through custom chain-of-thought prompting, but this is a fragile workaround, and being competent on SimpleToM without such guidance is thus an important measure of a model’s capabilities.
@@
We make our SimpleToM dataset and the full evaluation data for the analyzed models publicly available at https://huggingface.co/datasets/allenai/SimpleToM. This will allow researchers to build on top of our work in studying the neural ToM capabilities of LLMs in general, as well as further exploiting the diversity of situations covered by SimpleToM in exploring applied ToM.
@@ 2.1 Different types of information asymmetry
To expand beyond the classical false belief task, we seed the creation of SimpleToM with ten diverse scenarios where information asymmetry occurs naturally (Table 1). This is inspired by social psychology literature to cover asymmetries like manipulation, deception, secrecy, lying, and misleading behavior (Doherty, 2008), seen in real-world contexts like sales of “lemon” products, where items with hidden flaws are purchased due to a lack of information (Akerlof, 1978). These are under-examined in existing ToM tests. We further describe the scenarios with examples in Table 7 (Appendix E.1).
@@
Table 1: The ten broad scenarios used to seed the generation of stories in SimpleToM. Each scenario describes a type of information asymmetry that occurs naturally in the real world.
@@ 2.2 Simple story format without explicit percepts or mental states
We impose the constraint that Person X’s unawareness of the key information should be implicit (e.g., avoid explicit use of perception or mentalizing words such as “see”, “notice” or “belief”). This design encourages models to read between the lines and make commonsense inferences over the given situations and infer characters’ mental states in a more realistic manner, bringing us closer to realistic daily life use cases of ToM. (E.g., you cannot see through a Pringles can; you would not know about a cheating event if you were not present.)
@@
  • Unaware behavior: A likely next action by Person X given that they are unaware of the key information
  • Aware behavior: A likely next action by Person X if they were somehow aware of the key information after all (counterfactual)
@@ 2.3 Questions testing explicit and applied ToM
We use three types of questions (Figure 1) to probe a model’s grasp of each story, covering both explicit theory of mind (conceptual knowledge about others’ mental states; i.e., via (a) mental state question about information awareness) and applied theory of mind (the ability to use theory of mind in real-life contexts; i.e., via (b) behavior and (c) judgment questions) (Lee et al., 2024).
Mental state (MS) question about information awareness: We test ability of models to infer mental states, specifically information awareness, through a simple yes/no question (Is <Person X> likely to be aware that "<key information>"?) To infer whether a character is aware of something in SimpleToM stories, a model has to make implicit commonsense inferences about what the character can perceive or know in the given situation (including commonsense reasoning about physical objects, space, intent, goals of others, and so on).
Behavior prediction question: This question asks which of two possible actions the main subject (Person X) is likely to perform next. For instance, beyond answering that a person shopping for chips in the supermarket is unlikely to know that “the can of Pringles has moldy chips in it”, a model that successfully applies this inference for behavior prediction should also infer that a person who picked up such a can in the supermarket would likely “pay for the chips” rather than “report the moldy chips.” To answer these questions correctly, models need to reason over the situation to infer the mental state of character, and realize how the character’s lack of awareness of the key information would impact their likely next action.
Judgment question: The judgment question specifies that “correct” action was taken (rather than the incorrect one) and asks if this was a reasonable choice. Similar to the behavior question, the judgment question requires inferring and applying implicitly conveyed information like percepts, beliefs, desires, and knowledge. People’s mental states are an important factor to consider in making appropriate judgments of their behavior (Jara-Ettinger et al., 2016; Schein & Gray, 2018). For instance, buying a can of Pringles that has moldy chips in it is not a reasonable action if the person knows about the moldy chips. However, it is a perfectly reasonable (and expected) behavior if this piece of key information is not a part of the person’s mental state.
@@ 2. Step 2:
    For each scenario, using the seed story as example, prompt the LLM to suggest 10 diverse sets of entities compatible with an information asymmetry. (See prompt in E.5.)
@@ 3. Step 3:
    For each set of suggested entities, along with the seed story, prompt the LLM to write three new stories at different levels of “severity.” With each story, also generate likely next “unaware” and “aware” behaviors (see Section 2.2). Appendix  E.4 provides further details.
We went through two rounds of this process. First, we used GPT-4 and Claude-3-Opus [^1] See Table 6 for exact models used to generate a total of 1200 stories.22210 scenarios * 2 models * 10 entities per model * 3 severities * 2 models to generate stories After annotating and filtering this initial set (Section 3.2), we picked a new set of top-scoring seed stories and sourced 10 additional sets of entities from each of GPT-4o and Claude-3.5-Sonnet. We used these two newer models to generate stories for all 40 sets of entities, for a total of 2400 more stories. By using several generator models, varied entities and different seed stories, the resulting stories in SimpleToM have a wide range of information asymmetries instantiated in different real-world situations, effectively broadening neural ToM tests beyond traditional settings (Section 6). These contexts also allow for nuanced and implicit traits (e.g., buyers would avoid products with defects if they know about them).
@@ 3.2 Strict quality control on stories that goes into SimpleToM
We gather human annotations on each story (and unaware/aware next actions). We asked annotators four questions for each story, summarized in Figure 2. This process verifies that the key information in each story is something that Person X has false belief about. We also carefully verify that the next likely “unaware action” is appropriate if and only if Person X is unaware of the key information. We similarly verify the “aware action,” for the counterfactual situation where Person X is somehow aware of the key information. Appendix D provides further details about the crowdsourcing procedure, with instructions, examples and question templates.
Our annotators passed a rigorous qualification test (Appendix D.2) and met other high-standard requirements (Appendix D.3). Only stories for which all crowdworkers (3) judged all aspects to be valid were included in SimpleToM. [^3] See more details in Appendix D.4. This results in 1147 stories (out of the original 3600) in the final SimpleToM dataset. Table 8 (Appendix E.2) provides statistics and further details for SimpleToM.
@@ 4 Experimental setup
We evaluate SimpleToM on ten frontier LLMs from different sources and with different levels of capabilities: Llama-3.1-8B, Llama-3.1-405B, (Dubey et al., 2024) Claude-3-Haiku, Claude-3-Opus (Anthropic, 2024b), Claude-3.5-Sonnet,(Anthropic, 2024a) GPT-3.5, GPT-4, GPT-4o, o1-mini, and o1-preview (OpenAI, 2024; 2023) (refer to Appendix C Table 6 for more details). We use the most deterministic setting with a generation temperature of 0.444Except the o1 models which do not allow adjusting the temperature or system prompt.
@@
    How well can models (i) infer characters’ mental states, (ii) anticipate characters’ behavior and (iii) make appropriate judgments, requiring the use of ToM inferences?
@@
    How can the performance on the behavior and judgment questions improve with various levels intervention during inference, such as patching mental state inference in the prompt (MS remind), guiding the model via system prompt (SysP), and encouraging reasoning/thinking before answering (CoT)?
@@ 5.1 Frontier LLMs can infer mental states, but struggle to use it
The overall evaluation results on SimpleToM for the 10 models are summarized in Table 2, spanning the different question types (as detailed in Section 2.3). We analyze models’ performance for each type of question below. Note that these are binary questions where random performance is 50%.
@@
GPT-3.5 36.5 7.6 29.1 24.4
Claude-3-Haiku 87.2 23.6 16.7 42.5
o1-mini 87.8 44.8 27.0 53.2
GPT-4o 95.6 49.5 15.3 53.5
Llama-3.1-405B 97.8 58.2 10.0 55.4
Claude-3-Opus 98.3 64.4 9.6 57.4
GPT-4 96.6 63.0 19.5 59.7
Llama-3.1-8B 88.1 38.5 54.6 60.4
Claude-3.5-Sonnet 97.9 67.0 24.9 63.3
o1-preview 95.6 84.1 59.5 79.7
Table 2: Evaluation results for SimpleToM on the different question types. Models are generally proficient in explicit ToM questions (directly querying about “mental state”, i.e., information awareness) but this success does not transfer to applied ToM (“behavior” and “judgment” questions).
@@
Mental state (MS) question about information awareness: Our results (Table 2, “mental state” column) show that reasoning over implicit information in given situations to infer mental states is still challenging for models like GPT-3.5 (36.5% accuracy), while newer and/or bigger models like Claude-3-Haiku, o1-mini, and Llama-3.1-8B perform reasonably well (around 88%). In fact, all recent frontier models are proficient at inferring characters’ awareness in our dataset – GPT-4o, Llama-3.1-405B, Claude-3-Opus, GPT-4, Claude-3.5-Sonnet, and o1-preview all achieved accuracies of more than 95%. This result also confirms the quality of our dataset, in that characters’ mental states in SimpleToM stories are implicit but reasonably easy to infer, as designed.
@@
Behavior prediction: On behavior prediction questions (Table 2, “behavior” column), smaller and older models perform extremely poorly (with GPT-3.5 achieving only 7.6% accuracy and several other models scoring less than 40%). Even for the larger models, like Llama-3.1-405B, Claude-3.5-Sonnet, and GPT-4o, performance on the behavior prediction task is much worse than on the mental state task with at least a 30% performance drop. This large inconsistency suggests that while frontier LLMs may have the right conceptual knowledge/information about others’ mental states when directly asked, they struggle to apply this knowledge in real-life contexts to make predictions about characters’ behavior. Only the latest o1-preview model, with its built-in inference time reasoning tokens, [^5] The o1 reasoning tokens make these models more like the chain-of-thought prompted versions of the other models below, although without any custom prompt. See Appendix H for discussion of the number of output tokens used by the o1 and the other models when using chain-of-thought. manages a decent score on this question type (84.1%).
@@
Judgment of behavior: Our results (Table 2, “judgment” column) show that this additional, seemingly trivial, inference step (beyond the behavior prediction) makes the task much more difficult for all the models. Even the newest and largest models like Llama-3.1-405B, Claude-3.5-Sonnet, and GPT-4o, which all achieved accuracies of more than 95% on inferring characters’ mental state, consistently make inaccurate judgments of behavior, and their performance drop to far below random (with accuracies in the range of 10% to 24.9%). Llama-3.1-8B somehow manages around random performance (at 54.6%), as an outlier. Finally, o1-preview, the best-performing model on behavior prediction, also gets the best score on this judgment prediction task, but even so it reaches only 59.5%, far below its performance on the other questions. These results show that even the latest models lack the ability to effectively apply ToM reasoning in this kind of judgment task.
@@
Figure 3: Considering the sequence of first predicting mental state, then behavior and finally judgment, we can record a failure for the first mistake. The “fail at behavior prediction” and “fail at judgment” parts can be considered inconsistent predictions by the model, since it got the associated mental state (and behavior prediction) questions correct.
@@
Overall ToM reasoning: To fully reason about the judgment question, requires reasoning about the behavior prediction, which relies on understanding the mental state (information awareness) question. We can visualize this by recording the first failure on the three “mental state” $\rightarrow$ “behavior” $\rightarrow$ “judgment” questions. Figure 3 shows the distribution of such failures, showing the proportionally large cases failing at the behavior and judgment steps. For most models the green segment, representing full understanding of the stories in terms of ToM reasoning, is very small.

### v2

@@ Abstract
Large language models (LLMs) are increasingly tested for a “Theory of Mind” (ToM) — the ability to attribute mental states to oneself and others. Yet most evaluations stop at explicit belief attribution in classical toy stories or stylized tasks, leaving open the questions of whether LLMs can implicitly apply such knowledge to predict human behavior, or to judge an observed behavior, in diverse scenarios. We introduce SimpleToM, a benchmark that advances ToM evaluation along two novel axes. First, it probes multiple levels of ToM reasoning, from mental state inference (explicit ToM) to behavior prediction and judgment (applied ToM). Second, it situates these tasks in diverse, everyday scenarios — such as supermarkets, hospitals, schools, and offices — where information asymmetries naturally arise (e.g., hidden defects in grocery store items, incomplete information in provider–patient interactions, or restricted access to locked devices). SimpleToM contains concise stories (e.g., “The can of Pringles has moldy chips in it. Mary picks up the can in the supermarket and walks to the cashier.”), each with three questions that test different degrees of ToM reasoning, asking models to predict: (a) mental states (“Is Mary aware of the mold?”), (b) behaviors (“Will Mary pay for the chips or report the mold?”), and (c) judgments (“Mary paid for the chips. Was that reasonable?”). Experiments reveal a striking gap: state-of-the-art models often reliably infer mental state (a), but fail at applying knowledge about the mental state for secondary predictions, with performance dropping sharply for behavior prediction (b) and further for behavior judgment (c). This exposes a critical fragility in LLMs’ social reasoning in terms of what they know (explicit ToM) versus how well they can implicitly apply that knowledge for predictions (applied ToM). By uniting assessment of different levels of ToM reasoning with diverse, everyday scenarios, SimpleToM opens new opportunities for rigorously evaluating and diagnosing ToM abilities in LLMs, and reveals surprising, new insights about current model capabilities, guiding efforts toward future generations of models capable of robust social understanding.
@@ 1 Introduction
As LLMs are now regularly used as conversational agents, it is critical that they can reliably reason about other people’s beliefs. Without this, an LLM may provide disastrous responses, for example by failing to recognize emotional distress (Obradovich et al., 2024), treating a sarcastic comment as literal truth (Zhang et al., 2024), providing direct but inappropriate advice in sensitive situations (Hodson and Williamson, 2023; Kim et al., 2025), or blindly helping a user with malevolent intent (Shang et al., 2024). Performing such social reasoning is complex, involving attributing mental states to oneself and others, an ability widely known as Theory of Mind (ToM) (Premack and Woodruff, 1978). Specifically, ToM requires an LLM to reason over multiple, possibly conflicting views of the world simultaneously, making it fundamentally more challenging than typical multistep reasoning over a single worldview (e.g., factual, ontological, or arithmetic knowledge). This requires specific studies to ascertain how well LLMs perform on ToM, distinguishing between reasoning on the basis of the state of the world and reasoning on the basis of someone’s beliefs about the state of the world (Doherty, 2008), and such studies are becoming increasingly needed as LLM adoption in society grows.
@@
However, while ToM in humans has been extensively studied in psychology (e.g., Doherty, 2008; Baron-Cohen et al., 1985; Perner et al., 1987), studies of ToM reasoning in LLMs to date have been limited, largely relying on the classical Sally-Anne task or templated variants of it (Le et al., 2019; Nematzadeh et al., 2018; Wu et al., 2023; Xu et al., 2024). While informative, these studies have several shortcomings: (i) limited diversity in how information asymmetry arises (see related work in Section 6 for examples across existing datasets), (ii) explicit use of percept and mentalizing verbs like “sees” and “thinks” which serve as trigger words for models to realize that these are important aspects, removing the need for implicit commonsense inferences about relevant percepts or beliefs, and (iii) limited exploration of applied ToM, such as the judgment of behavior which requires implicit reasoning about mental state.
@@
Our goal is to go beyond assessing just mental state inference (“What does X believe?”), to also assess how well models can predict others’ behavior based on that understanding (“What will X do?”), and make judgments of appropriateness of that behavior (“Did X act appropriately?”), as well as to unite assessment of these different levels of ToM reasoning with diverse, everyday scenarios – e.g., supermarkets, schools, offices – where information asymmetries naturally arise. Our approach is to construct and evaluate using a new dataset, SimpleToM. Each story in SimpleToM is paired with three types of questions targeting these abilities (Figure 1), with a total of 1147 stories and 3441 questions in daily life settings. Our results are surprising, revealing a significant gap in model performance between explicit and applied ToM questions (Hutchins et al., 2016; Lee et al., 2024), even though the underlying scenarios are identical. We find that frontier models perform well on explicit ToM questions (directly querying for information about “mental state”, i.e., information awareness). However, this success does not extend to applied ToM (“behavior” and “judgment” questions), even in strong models like GPT-5 and o1-preview. Performance can also vary wildly across scenarios, highlighting the importance of SimpleToM covering diverse scenarios beyond those in classical ToM tests, for rigorously evaluating and diagnosing ToM abilities in LLMs. Overall, the results show that frontier models still lack the ability to independently and reliably apply ToM skills in tasks such as anticipating others’ behavior and making judgments, calling for caution when using them in social applications (see discussion of example applications in Appendix A).
Our contributions and findings are as follows:
@@
  • We introduce SimpleToM, a dataset for testing the core abilities of LLMs in both explicit and applied ToM.
@@
  • We find that current frontier models have decoupled capabilities between predicting someone’s information awareness in a situation (explicit ToM, which they excel at), and utilizing it to predict and judge someone’s behavior (applied ToM, which they perform poorly at).
  • We show that LLMs’ ToM performance can vary wildly across scenarios, highlighting the importance of using diverse scenarios for diagnosing ToM abilities in LLMs.
@@
We make our SimpleToM dataset [^1] https://huggingface.co/datasets/allenai/SimpleToM and code [^2] https://github.com/yulinggu-cs/SimpleToM publicly available. This will allow researchers to build on top of our work in studying the neural ToM capabilities of LLMs in general, as well as to further exploit the different levels of ToM reasoning and diversity of situations covered by SimpleToM.
@@ 2.1 Different types of information asymmetry
To expand beyond the classical false belief, or Sally-Anne task, we seed the creation of SimpleToM with ten diverse scenarios where information asymmetry occurs naturally in everyday settings such as in supermarkets, schools, and offices (Table 1). This is inspired by social psychology literature to cover asymmetries like manipulation, deception, secrecy, lying, and misleading behavior (Doherty, 2008) seen in real-world contexts like sales of “lemon” products, where items with hidden flaws are purchased due to a lack of information (Akerlof, 1978). These are under-examined in existing ToM tests. We further describe the scenarios with examples in Table 8 (Appendix I.1).
Table 1: The ten broad scenarios used to seed the generation of stories in SimpleToM. Each scenario describes a type of information asymmetry that occurs naturally in everyday setting.
@@ 2.2 Simple story format without explicit percepts or mental states
We impose the constraint that Person X’s unawareness of the key information should be implicit (e.g., avoid explicit use of perception or mentalizing words such as “see”, “notice” or “believe”). This design encourages models to read between the lines and make commonsense inferences over the given situations and infer characters’ mental states in a more realistic manner, bringing us closer to realistic daily life use cases of ToM. (E.g., you cannot see through a Pringles can; you would not know about a cheating event if you were not present.)
@@
  • Unaware behavior: A likely next action by Person X given that they are unaware of the key information.
  • Aware behavior: A likely next action by Person X if they were somehow aware of the key information after all (a counterfactual).
@@ 2.3 Questions testing explicit and applied ToM
We use three types of questions (Figure 1) to probe a model’s grasp of each story, covering both explicit theory of mind (conceptual knowledge about others’ mental states; i.e., via (a) mental state questions about information awareness) and applied theory of mind (the ability to use theory of mind in downstream tasks i.e., via (b) behavior and (c) judgment prediction questions) (Hutchins et al., 2016; Lee et al., 2024).
Mental state (MS) question about information awareness: We test ability of models to infer mental states, specifically information awareness, through a simple yes/no question (Is <Person X> likely to be aware that "<key information>"?). To infer whether a character is aware of something in SimpleToM stories, a model has to make implicit commonsense inferences about what the character can perceive or know in the given situation (including commonsense reasoning about physical objects, space, intent, goals of others, and so on).
Behavior prediction question: This question asks which of two possible actions the main subject (Person X) is likely to perform next. For instance, beyond answering that a person shopping for chips in the supermarket is unlikely to know that “the can of Pringles has moldy chips in it”, a model that successfully applies this inference for behavior prediction should also infer that a person who picked up such a can in the supermarket would likely “pay for the chips” rather than “report the moldy chips.” To answer these questions correctly, models need to implicitly reason over the situation to infer the mental state of character(s), and realize how the character’s lack of awareness of the key information would impact their likely next action.
Judgment question: The judgment question specifies that the “correct” action was taken (rather than the incorrect one) and asks if this was a reasonable choice. As the inference graph in Figure 1 illustrates, the judgment question goes beyond behavior prediction as it requires two levels of implicit reasoning, first implicitly predicting the behavior of Person X, which itself relies on implicitly understanding their mental state. People’s mental states are an important factor to consider in making appropriate judgments of their behavior (Jara-Ettinger et al., 2016; Schein and Gray, 2018). For instance, buying a can of Pringles that has moldy chips in it is not a reasonable action if the person knows about the moldy chips. However, it is a perfectly reasonable (and expected) behavior if this piece of key information is not a part of the person’s mental state.
@@ 2. Step 2:
    For each scenario, using the seed story as example, prompt the LLM to suggest 10 diverse sets of entities compatible with an information asymmetry. (See prompt in Appendix I.5.)
@@ 3. Step 3:
    For each set of suggested entities, along with the seed story, prompt the LLM to write three new stories at different levels of “severity.” With each story, also generate likely next “unaware” and “aware” behaviors (see Section 2.2). Appendix  I.4 provides further details.
We went through two rounds of this process. First, we used GPT-4 and Claude-3-Opus [^3] See Table 4 for exact models used. to generate a total of 1200 stories.44410 scenarios * 2 models * 10 entities per model * 3 severities * 2 models to generate stories After annotating and filtering this initial set (Section 3.2), we picked a new set of top-scoring seed stories and sourced 10 additional sets of entities from each of GPT-4o and Claude-3.5-Sonnet. We used these two newer models to generate stories for all 40 sets of entities, for a total of 2400 more stories. By using several generator models, varied entities and different seed stories, the resulting stories in SimpleToM have a wide range of information asymmetries instantiated in different everyday situations, effectively broadening neural ToM tests beyond traditional settings (Section 6). These contexts also allow for nuanced and implicit traits (e.g., buyers would avoid products with defects if they know about them).
@@ 3.2 Strict quality control on stories that goes into SimpleToM
We gather human annotations on each story (and unaware/aware next actions). We asked annotators four questions for each story, summarized in Figure 2. This process verifies that the key information in each story is something that Person X has a false belief about. We also carefully verify that the next likely “unaware action” is appropriate if and only if Person X is unaware of the key information. We similarly verify the “aware action” for the counterfactual situation where Person X is somehow aware of the key information. Appendix G provides further details about the crowdsourcing procedure, with instructions, examples and question templates.
Our annotators passed a rigorous qualification test (Appendix G.2) and met other high-standard requirements (Appendix G.3). Only stories for which all crowdworkers (3) judged all aspects to be valid were included in SimpleToM. [^5] See more details in Appendix G.4. This results in 1147 stories (out of the original 3600) in the final SimpleToM dataset. Table 9 (Appendix I.2) provides statistics and further details for SimpleToM.
@@ 4 Experimental setup
We evaluate SimpleToM on 21 LLMs from different sources and with different levels of capabilities, ranging from smaller open-weights models to close-sourced frontier models: Llama-3.1-8B, Llama-3.1-405B, Llama-3.2-1B, Llama-3.2-3B (Dubey et al., 2024), Qwen-2.5-7B, Qwen-2.5-14B (Qwen et al., 2025), Ministral-8B (Mistral, 2024), Claude-3-Haiku, Claude-3-Opus (Anthropic, 2024b), Claude-3.5-Sonnet (Anthropic, 2024a), DeepSeek-R1 (DeepSeek-AI et al., 2025), GPT-3.5, GPT-4, GPT-4.5-preview, GPT-5, GPT-4o, GPT-4o-mini, o1-mini, o3-mini, o1 and o1-preview (OpenAI, 2023; 2024) (refer to Appendix E Table 4 for more details). Where possible, we use the most deterministic setting with a generation temperature of 0.666For test-time reasoning models (o1, o3, GPT-5 and DeepSeek-R1 models) we use the recommended temperature settings.
@@
    How well can models (a) infer characters’ mental states, (b) anticipate characters’ behavior and (c) make appropriate judgments, requiring the use of ToM inferences?
@@
    How much can we close the gap between models’ performance on explicit and applied ToM using inference-time interventions?
@@ 5.1 Frontier LLMs can infer mental states, but struggle to use it
The overall evaluation results on SimpleToM for the 21 models are summarized in Table 2, spanning the different question types (as detailed in Section 2.3). We analyze models’ performance for each type of question below. Note that these are binary questions where random performance is 50%. For each score we also report the 95% confidence interval. [^7] Using Wald interval for binomial distribution based on 1147 samples in each category.
Table 2: Evaluation results for SimpleToM on the different question types. State-of-the-art models are generally proficient in explicit ToM questions (directly querying about “mental state”, i.e., information awareness) but this success does not transfer to applied ToM (“behavior” and “judgment” questions). Each score is annotated with a 95% confidence interval.
@@
GPT-3.5 36.5±2.8 7.6±1.5 29.1±2.6 24.4±1.4
Qwen-2.5-7B 76.4±2.5 25.3±2.5 23.4±2.5 41.7±1.6
Claude-3-Haiku 87.2±1.9 23.6±2.5 16.7±2.2 42.5±1.7
Ministral-8B 62.5±2.8 26.9±2.6 50.0±2.9 46.5±1.7
Qwen-2.5-14B 93.5±1.4 35.3±2.8 15.5±2.1 48.1±1.7
Llama-3.2-3B 62.1±2.8 32.9±2.7 49.8±2.9 48.2±1.7
GPT-4o-mini 93.0±1.5 40.6±2.8 22.3±2.4 52.0±1.7
GPT-4o 95.6±1.2 49.5±2.9 15.3±2.1 53.5±1.7
Llama-3.2-1B 91.8±1.6 22.1±2.4 49.3±2.9 54.4±1.7
Llama-3.1-405B 97.8±0.8 58.2±2.9 10.0±1.7 55.4±1.7
Claude-3-Opus 98.3±0.7 64.4±2.8 9.6±1.7 57.4±1.7
GPT-4 96.6±1.0 63.0±2.8 19.5±2.3 59.7±1.6
Llama-3.1-8B 88.1±1.9 38.5±2.8 54.6±2.9 60.4±1.6
Claude-3.5-Sonnet 97.9±0.8 67.0±2.7 24.9±2.5 63.3±1.6
GPT-4.5-preview 97.0±1.0 67.8±2.7 26.7±2.6 63.8±1.6
Reasoning models (with internal chain of thought)
o1-mini 87.8±1.9 44.8±2.9 27.0±2.6 53.2±1.7
o1 98.6±0.7 58.8±2.8 32.5±2.7 63.3±1.6
o1 (high reasoning effort) 98.7±0.7 60.2±2.8 33.3±2.7 64.1±1.6
o3-mini 85.4±2.0 66.8±2.7 41.3±2.8 64.5±1.6
GPT-5 98.5±0.7 64.4±2.8 40.0±2.8 67.7±1.6
DeepSeek-R1 97.3±0.9 73.8±2.5 65.8±2.7 79.0±1.4
o1-preview 95.6±1.2 84.1±2.1 59.5±2.8 79.7±1.3
@@
Mental state (MS) question about information awareness: Our results (Table 2, “mental state” column) show that reasoning over implicit information in given situations to infer mental states is still challenging for models like GPT-3.5 (36.5% accuracy), while newer and/or bigger models like Claude-3-Haiku, Llama-3.1-8B, o1-mini perform reasonably well (around 88%). In fact, all recent frontier models are proficient at inferring characters’ awareness in our dataset e.g., models like GPT-4o, Llama-3.1-405B, Claude-3-Opus, GPT-4, GPT-5, Claude-3.5-Sonnet, o1-preview, o1, DeepSeek-R1 all achieved accuracies of more than 95%. This result also confirms the quality of our dataset, in that characters’ mental states in SimpleToM stories are implicit but reasonably easy to infer, as designed.
@@
Behavior prediction: On behavior prediction questions (Table 2, “behavior” column), smaller and older models perform extremely poorly (with GPT-3.5 achieving only 7.6% accuracy, and models like Claude-3-Haiku and Llama-3.1-8B scoring less than 40%). Even for the larger models, like Llama-3.1-405B, Claude-3.5-Sonnet, GPT-4, GPT-5 and GPT-4o, performance on the behavior prediction task is much worse than on the mental state task with at least a 30% performance drop. [^8] See Appendix 10 for how humans demonstrate greater consistency across the question types. This large inconsistency suggests that while frontier LLMs may have the right conceptual knowledge/information about others’ mental states when directly asked, they struggle to apply this knowledge in everyday scenarios to make downstream predictions about characters’ behavior. Only the o1-preview model, with its built-in inference time reasoning tokens, [^9] The o1 reasoning tokens make these models more like the chain-of-thought prompted versions of the non-reasoning models, although without any custom prompt. See Appendix L for discussion of the number of output tokens used by the o1 and the other models when using test-time reasoning tokens. manages a decent score of more than 80% on this question type (84.1%).
@@
Judgment of behavior: Our results (Table 2, “judgment” column) show that this additional inference step (making judgment of the characters’ behavior beyond just behavior prediction) makes the task much more difficult for all the models. Even the newer and larger models like Llama-3.1-405B, Claude-3.5-Sonnet, and GPT-4o, which all achieved accuracies of more than 95% on inferring characters’ mental state, consistently make inaccurate judgments of behavior, and their performances drop to far below random (with accuracies in the range of 10% to 24.9%). A subset of models (Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, Ministral-8B) manages performance near random chance level, while performing relatively worse on behavior prediction, presumably because these models are failing to make a meaningful judgment beyond random guessing. Even the best performing model on judgment prediction, DeepSeek-R1, reaching 65.8%, is significantly below its performance on the other questions, illustrating a clear explicit vs. applied ToM performance gap.
@@
Figure 3: Considering the sequence of first predicting mental state, then behavior and finally judgment, we record a failure for the mistake that occurs first. The “fail at behavior prediction” and “fail at judgment” demonstrate inconsistent predictions by the model, since the model got the associated mental state (and behavior prediction) question(s) correct.
@@
Summary: To fully reason about the judgment question requires reasoning about the behavior prediction, which in turn relies on understanding the mental state (information awareness) question. We can visualize this by recording the first failure on the three “mental state” $\rightarrow$ “behavior” $\rightarrow$ “judgment” questions. Figure 3 shows the distribution of such failures, showing the large proportion of cases failing at the behavior and judgment steps (applied ToM). For most models the “All Correct” segment, representing full understanding of the 2-sentence stories in SimpleToM in terms of ToM reasoning, is very small.